SHORT ANSWER. Write the word or phrase that best completes each statement or answers the ques tion. For a)- db identify each probability as Marginal, Conditional, or Joint. Then find the probability as a de rounded to the nearest tenth place. 5) The managers of a corporation were surveyed to desermine the background that keads to_ a successful manager. Eachmanager was rated as being either a good tale or poor manager by his/her boss The managers ed ucational background was aso nobed. The data appear below Educational Background Manager Rating 1. S. Degree Good Fair Poor Totals 17 28 18 34 a) What is the probability that a manager has a College Degree b) Given that a manager is only a fair manager, what is the probability that this manage has no college background? c) What is the probability that a someone is a poor manager, given the he or she has a college degree? d) What is the probability that a manager has a poor rating and has a college degree?
Answer #1

Ans;

HS some College Master/phd Total
Good 5 6 24 4 39
Fair 8 18 44 17 87
Poor 3 2 1 28 34
Total 16 26 69 49 160
Joint probability distribution:
HS some College Master/phd Total
Good 0.03125 0.0375 0.15 0.025 0.24375
Fair 0.05 0.1125 0.275 0.10625 0.54375
Poor 0.01875 0.0125 0.00625 0.175 0.2125
Total 0.1 0.1625 0.43125 0.30625 1

a)P(college degree)=69/160=0.43125

b)No college background means that have HS degree

P(HS degree/fair)=P(HS degree and fair)/P(fair)=(8/160)/(87/160)=8/87=0.0920

c)P(poor/college degree)=P(poor and college degree)/P(college degree)

=(1/160)/(69/160)=1/69=0.0145

d)P(poor and college degree)=1/160=0.00625